# Robust $\mathcal{H}_{\infty}$ control under stochastic requirements: minimizing conditional value-at-risk instead of worst-case performance

### v2

@@ Abstract
Conventional robust $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ control minimizes the worst-case performance, often leading to a conservative design driven by very rare parametric configurations. To reduce this conservatism while taking advantage of the stochastic properties of Monte-Carlo sampling and its compatibility with parallel computing, we introduce an alternative paradigm that optimizes the controller with respect to a stochastic criterion, namely the conditional value at risk. We illustrate the potential of this approach on a realistic satellite benchmark, showing that it can significantly improve overall performance by tolerating some degradation in very rare worst-case scenarios.
I Introduction
Robust structured $\mathcal{H}_{\infty}$ control [5] traditionally aims at finding a controller that minimizes the worst-case performance across all possible parametric configurations:
@@ I Introduction
where $T_{zw}(\mathrm{s},k,\delta)$ is a transfer function between input $w$ and output $z$, $k$ is a vector of decision variables defined over an admissible space $\mathcal{K}$ and describing a controller $K$ of chosen structure (fixed-order, PID, etc.), and $\delta$ is a vector of uncertain parameters. In this formulation, the space $\mathcal{D}_{\delta}$ of uncertain parameters is deterministic, and is generally the hypercube $1,1]^{m}$ after normalization of the $m$ uncertain parameters. The case of nonlinear constraint restricting this parametric space was also addressed in [15]. When there are multiple control objectives, it is sought to solve:
@@
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm. The channel $w_{1}\rightarrow z_{1}$ corresponds to a soft requirement to be minimized, while the other channels $w_{i}\rightarrow z_{i}$ ($i>1$) are hard requirements, that must be inferior to 1 (once normalized) but not necessarily minimized. As this formulation does not involve any notion of probability, the worst-case configuration is often extremely unlikely, leading to a conservative design with regard to industrial requirements that are generally defined in a stochastic sense and relying on a probabilistic description of the uncertain parameters (e.g. gaussian, uniform, or other distributions).
@@
When considering deterministic parameters $\delta\in[-1,1]^{m}$, the state-of-art for robust $\mathcal{H}_{\infty}$ control is the nonsmooth optimization algorithm of [2], implemented in MATLAB’s function systune, that solves the inner maximization problem of (1) using a local search enabled by the computation of Clarke subdifferentials [11] of the $\mathcal{H}_{\infty}$ norm [9]. Identified worst-case configurations are iteratively added to an active set $\mathcal{D}_{a}\subset\mathcal{D}$ on which the outer minimization problem is solved based on the algorithm of [4]. Then, global robustness verification can be done for example with Monte-Carlo sampling [27] or with $\mu$-analysis [12]. The former provides statistical guarantees based on the worst-case sample, while the latter, generally based on branch-and-bound algorithms [21], provides a formal guarantee of detecting the global worst-case configuration. However, as noted above, this worst-case scenario can be extremely rare, which motivated the development of probabilistic $\mu$-analysis [8]. Nevertheless, as most global optimization methods, $\mu$-based techniques exhibit limited scalability with the number of parameters; in fact, they attempt to solve a NP-hard problem [28]. Furthermore, they are not oriented toward robust controller synthesis, which is the primary concern of the present work.
@@
This paper proposes an alternative paradigm to perform robust $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ controller synthesis, with two objectives: (i) to optimize a robust controller in the sense of stochastic requirements, thereby ensuring a better match between the control synthesis framework and the stochastic model which better reflects the true underlying uncertainty-generating process than the deterministic formulation, thus reducing conservatism, and (ii) to leverage Monte-Carlo methods, which are commonly used in stochastic programming [23], come with stochastic guarantees, and naturally exploit parallel computing.
@@
For this, we rely on the conditional value at risk (CVAR), a common concept in the field of finances to quantify potential extreme losses in a portfolio. The main idea of this approach is to consider $T_{zw}(\mathrm{s},k,\delta)||$ as a random variable defined over a probabilistic space rather than a deterministic function defined over a hypercube. This formulation is introduced in Section II, and adapted in Section III to be more suitable for optimization. Section IV-A proposes a solution method and identifies open challenges. Finally, Section V illustrates this method on the robust control of a mechanical system.
We note that the use of CVAR in control problems is not new; see, for example, [13, 16, 29]. However, to the best of the authors’ knowledge, its application to the robust $\mathcal{H}_{\infty}$ control problem, traditionally treated via the worst-case approach described above, has not been previously discussed.
II Problem statement
@@
Assuming a bounded support is common in control engineering, as it is necessary in the traditional deterministic framework to define a worst case (as in Eq. (1)). In this paper, it is necessary to realistically make Assumption 2 below. Without loss of generality, we consider that $\mathcal{D}_{\delta}\subset[-1,1]^{m}$ and that $\delta=0$ corresponds to the nominal configuration.
@@
Assumption 2
Given $k\in\mathcal{K}$, $L_{zw}(k,\delta)$ satisfies Assumption 2 if $T_{zw}(\mathrm{s},k,\delta)$ is internally stable for all $\delta\in\mathcal{D}_{\delta}$.
Assumption 2 allows the loss function to be well defined for any $\delta$, and thus to define the expected value in the following. This assumption might be seen as restrictive; but in practice, it can anyway be desirable to guarantee stability even in rare worst-case scenarios, with the following stochastic formulation relaxing only the performance requirements measured with the $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ norms.
Assumption 3
Given $k\in\mathcal{K}$, $L_{zw}(k,\delta)$ satisfies Assumption 3 when its distribution function is continuous.
Assumption 3 is adopted here following the work of [20] to properly introduce Definition 1.
Definition 1 ($\beta$-VAR, $\beta$-CVAR)
Given the loss function $L_{zw}(k,\delta)$ satisfying Assumptions 1, 2 and 3, the $\beta$-Value at risk ($\beta$-VAR), noted $\text{VAR}_{\beta}^{zw}(k)$, is:
@@
where $\Psi^{zw}(k,\alpha)$ is the probability of $L_{zw}(k,\delta)$ not exceeding a threshold $\alpha$, i.e. the cumulative distribution function. The $\beta$-Conditional value at risk ($\beta$-CVAR) is:
$\textrm{CVAR}_{\beta}^{zw}(k)=\frac{1}{1-\beta}\int_{L_{zw}(k,\delta)\geq\textrm{VAR}_{\beta}(k)}L_{zw}(k,\delta)p(\delta)d\delta\;.$ (5)
In other words, when looking at the distribution of probability of the loss function $L_{zw}(k,\delta)$, $\text{VAR}_{\beta}^{zw}(k)$ separates the $\beta$ most favorable cases from the $(1-\beta)$ worst cases, while $\textrm{CVAR}_{\beta}^{zw}(k)$ represents the conditional expected value of the loss associated with the worst cases $\delta$ relative to that loss being greater than $\text{VAR}_{\beta}^{zw}(k)$. This is schematically represented in Fig. 2.
Figure 2: Probability density $p^{zw}_{k}(\alpha)$ of the loss function $T_{zw}(\mathrm{s},k,\delta)||$ with $\beta$-VAR, $\beta$-CVAR and deterministic worst case (only defined under Assumption 1).
@@
which replaces problem (2), with the main differences: (i) the parameters $\delta$ are defined as random variables rather than deterministic parameters, over the same bounded support (Assumption 1), (ii) instead of minimizing the worst case without considering any probability, we seek to minimize the conditional expected value of the ($1-\beta$) worst cases.
III Computation of VAR and CVAR
III-A Formulation as an optimization problem
Definition 1 does not allow for a straightforward computation of the VAR and CVAR. Instead, one typically relies on the following result of Rockafellar, in which both are characterized as solutions of a convex optimization problem.
Proposition 1 (Expression of $\beta$-VAR and $\beta$-CVAR as solutions of a convex minimization problem [20])
Under Assumptions 1, 2 and 3, let us define the function
@@ III-A Formulation as an optimization problem
Corollary 1 (Optimal controller for $\beta$-CVAR minimization)
@@
III-B Sample average approximation (SAA)
The function $F^{zw}_{\beta}(k,\alpha)$ can be approximated by sampling NN i.i.d. realizations $\tilde{\delta}^{1}$, $\tilde{\delta}^{2}$, … $\tilde{\delta}^{N}$ from the density $p$ and defining the sample average approximation (SAA) [14]
@@ III-B Sample average approximation (SAA)
where $A(k,\delta)$ is the dynamics matrix of the state-space representation of the closed loop, and $\lambda^{\mathcal{R}}_{\max}(\cdot)$ denotes the spectral abscissa. The constraint on the spectral abscissa ensures that Assumption 2 remains satisfied over the whole support $\mathcal{D}_{\delta}$ and not only over the samples $\delta^{(i)}$. How this constraint can be enforced is discussed in Section IV-A.
III-C Interval of confidence
To obtain an accurate solution to the original problem, it is necessary to choose NN sufficiently large. [10, Theorem 3] characterizes the convergence rate and derives the number of samples required to guarantee a prescribed accuracy level with a given confidence. However, this result also depends on the regularity of the uncertain loss function, which is not known a priori. In practice, thanks to the expression of VAR and CVAR as solutions of an optimization problem (Proposition 1), the procedure proposed in [17, Section 3.2] allows to assess the gap between the estimated $\beta$-CVAR:
@@ III-C Interval of confidence
IV Optimization
IV-A Optimization procedure
@@
Initialization: Active configurations $\mathcal{D}_{\delta}^{a}=\{0\}$, number of samples NN, tolerance $\epsilon$.
Step 1: Find $\hat{k}_{0}$ that robustly stabilizes over the whole support $\mathcal{D}_{\delta}$.
Step 2: Form the SAAs with NN samples, initialize $k=\hat{k}_{0}$ and solve the SAA problem (12) where robust stability over the whole support:
@@
If $\lambda^{\mathcal{R}}_{\max}(A(k,\delta^{*}))>0$, add $\delta^{*}$ to $\mathcal{D}_{\delta}^{a}$ and go back to step 2. Otherwise, go to step 4.
@@
Step 4: Evaluate the CVARs of problem (12), with sufficiently tight intervals of confidence. If they are different from those obtained in step 2 by more than a factor $\epsilon$, increase NN and go back to step 2. Otherwise, terminate.
The stability over the whole support $\mathcal{D}_{\delta}$ is addressed through the active configurations approach of [2]; it consists of enforcing stability of configurations belonging to a subset $\mathcal{D}_{\delta}^{a}\subset\mathcal{D}_{\delta}$, initialized with the nominal configuration, and updating this subset through (deterministic) worst-case searches. This procedure is necessary even when the initial controller stabilizes the system over the whole support, because this property may not be maintained when solving the SAA problem (12).
@@
Step 1 allows to provide an initial controller with robust stability over the whole support, as required by Assumption 2. It can be performed with [2], which relies on the algorithm of [4] and the active configurations approach mentioned above; the pure stabilization problem is addressed with the shifted $\mathcal{H}_{\infty}$ norm [9]. One possible, albeit not mandatory, approach is to solve the deterministic problem (2) to obtain an initial controller which also guarantees robust performance with respect to the worst case. The optimization of the SAA approximation is performed in in step 2; it can be addressed with deterministic nonsmooth optimization. The worst-case search for unstable configurations is performed in step 3, as previously discussed. Finally, step 4 evaluates the estimated CVARs obtained with the optimized controller and their intervals of confidence [17] to validate or invalidate the controller.
IV-B Open challenges
@@
While we have proposed a method to address the CVAR optimization, there are still open challenges to address the formulated problem in a more systematic and efficient way.
@@
Future research should focus on developing suitable nonsmooth algorithms to efficiently address step 2 of Algorithm 1, with convergence certificates and able to start from any initial controller (i.e. not requiring step 1). Of particular interest is the work of [18], which proposes an algorithm that tolerates inexact computation of the gradient or of the function itself. This algorithm is specifically designed for lower-$C^{1}$ functions, a class to which the SAA $\widetilde{F}^{zw}_{\beta}(k,\alpha)$ can easily be shown to belong under Assumptions 1, 2 and 3, since this property holds for $T_{zw}(s,k,\delta)||_{\infty}$ as a function of $k$ as long as $T_{zw}(s,k,\delta)$ is internally stable [2, Proposition 2]. In particular, [18] illustrates their method on the $\mathcal{H}_{\infty}$ norm where the exactness of the computation, based on a dichotomy involving the bounded real lemma, can be relaxed. This property is particularly desirable in our SAA where we have to evaluate many times the $\mathcal{H}_{\infty}$ norm, and where our SAA is itself an approximation.
@@
V Application
V-A Benchmark
We consider the benchmark of [15], namely the attitude control of a flexible spacecraft $G(s,\delta)$ to follow a reference $r$. The closed loop is represented in Fig. 3, where $d(s)$ represents actuator dynamics. The control input $u$ of the plant is the torque applied to the satellite and the output vector contains the 3 angle measurements polluted by some noise $n$. We note $k\in\mathcal{K}$ the decision variables describing the controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, each one taking the measured angle and rate as inputs and generating a control torque. The closed-loop system has 50 states, 43 uncertain parameters $\delta_{i}$, and 72 decision parameters $k_{i}$.

### v3

# Robust $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ control under stochastic requirements: minimizing conditional value-at-risk instead of worst-case performance

@@ Abstract
Conventional robust $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ control minimizes the worst-case performance, often leading to a conservative design driven by very rare parametric configurations. To reduce this conservatism while taking advantage of the stochastic properties of Monte Carlo sampling and its compatibility with parallel computing, we introduce an alternative paradigm that optimizes the controller with respect to a stochastic criterion, namely the conditional value at risk. We present the problem formulation and discuss several open challenges toward a general synthesis framework. The potential of this approach is illustrated on a mechanical system, where it significantly improves overall performance by tolerating some degradation in very rare worst-case scenarios.
{IEEEkeywords}
Robust $\mathcal{H}_{\infty}$ control, stochastic optimization, risk-aware control, uncertain linear systems.
1 Introduction
@@
Robust structured $\mathcal{H}_{\infty}$ control [5] traditionally aims at finding a controller that minimizes the worst-case performance across all possible parametric configurations. Denoting as $T_{zw}(\mathrm{s},k,\delta)$ a transfer function from input $w$ to output $z$, which depends on the vectors $k$ of decision variables and $\delta$ of uncertain parameters, this formulation reads
@@
Here, $\mathcal{K}$ denotes the admissible space, describing the controllers with a chosen structure (fixed-order, PID, etc.), and $\mathcal{D}_{\delta}$ is generally the hypercube $1,1]^{m}$ after normalization of the $m$ uncertain parameters. The case of a nonlinear constraint restricting $\mathcal{D}_{\delta}$ was also addressed in [16]. When there are multiple control objectives, it is sought to solve:
[…]
@@
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm. The channel $w_{1}\rightarrow z_{1}$ corresponds to a soft requirement to be minimized, while the other channels $w_{i}\rightarrow z_{i}$ ($i>1$) are hard requirements, that must be less than 1 (once normalized) but not necessarily minimized. As this formulation does not involve any notion of probability, the worst-case configuration is often extremely unlikely, leading to a conservative design with regard to industrial requirements that are generally defined in a stochastic sense and relying on a probabilistic description of the uncertain parameters (e.g. gaussian, uniform, or other distributions).
@@
When considering deterministic parameters $\delta\in[-1,1]^{m}$, the state of the art for robust $\mathcal{H}_{\infty}$ control is the nonsmooth optimization algorithm of [2], implemented in MATLAB’s function systune [25], that solves the inner maximization problem of (1) using a local search enabled by the computation of Clarke subdifferentials [11] of the $\mathcal{H}_{\infty}$ norm [9]. Identified worst-case configurations are iteratively added to an active set $\mathcal{D}_{a}\subset\mathcal{D}_{\delta}$ on which the outer minimization problem is solved based on the algorithm of [4]. Then, global robustness verification can be done for example with Monte Carlo sampling [28] or with $\mu$-analysis [13]. The former provides statistical guarantees based on the worst observed sample, while the latter, generally based on branch-and-bound algorithms [21], provides a formal guarantee of detecting the global worst-case configuration. However, as noted above, this worst-case scenario can be extremely rare, which motivated the development of probabilistic $\mu$-analysis [8]. Nevertheless, $\mu$-based techniques exhibit limited scalability with the number of parameters, since they rely on bounding the structured singular value, whose exact computation is NP-hard [29]. Furthermore, they are not oriented toward robust controller synthesis, which is the primary concern of the present work.
@@
This paper proposes an alternative paradigm for robust $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ controller synthesis with two main objectives. The first is to optimize a robust controller in the sense of stochastic requirements, better aligning the synthesis framework with the underlying uncertainty-generating process and thereby reducing conservatism. The second is to leverage Monte Carlo methods, which are commonly used in stochastic programming [23], provide stochastic guarantees, and naturally exploit parallel computing.
@@
For this, we rely on the conditional value at risk (CVAR), a common concept in the field of finance to quantify potential extreme losses in a portfolio. The rationale for employing CVAR over other risk measures, such as the quantile, is twofold. First, it captures both the quantile threshold and the severity of losses beyond it, providing a more complete measure of risk. Second, it is particularly suitable for optimization, as established in the literature [20]. This formulation is introduced in Section 2, and adapted in Section 3 to be more suitable for optimization. Section 4.1 presents a first step toward a solution method and identifies several open challenges that remain to be addressed. Finally, Section 5 illustrates this method on the robust control of a mechanical system.
2 Problem statement
@@
Assuming a bounded support is common in control engineering, as it is necessary in the traditional deterministic framework to define a worst case (as in Eq. (1)). In this paper, it is necessary to realistically make Assumption 2.2 below. Without loss of generality, we consider that $\mathcal{D}_{\delta}\subset[-1,1]^{m}$ and that $\delta=0$ corresponds to the nominal configuration.
@@
Assumption 2.2
Given $k\in\mathcal{K}$, the loss function $L_{zw}(k,\delta)$ is said to satisfy Assumption 2.2 when $T_{zw}(\mathrm{s},k,\delta)$ is internally stable for all $\delta\in\mathcal{D}_{\delta}$.
Assumption 2.2 ensures that the loss function is well-defined for any $\delta$, and thus allows us to define the expected value in the following. This assumption might be seen as restrictive; but in practice, it can anyway be desirable to guarantee stability even in rare worst-case scenarios, with the following stochastic formulation relaxing only the performance requirements measured with the $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ norms.
Assumption 2.3
Given $k\in\mathcal{K}$, the loss function $L_{zw}(k,\delta)$ is said to satisfy Assumption 2.3 when its distribution function $\Psi^{zw}(k,\alpha)$ is continuous.
Assumption 2.3 is adopted here following the work of [20] to properly introduce Definition 2.4.
Definition 2.4
($\beta$-VAR, $\beta$-CVAR)
Given a loss function $L_{zw}(k,\delta)$ satisfying Assumptions 2.1, 2.2 and 2.3, and $\beta\in[0,1)$, the $\beta$-Value at risk ($\beta$-VAR), noted $\text{VAR}_{\beta}^{zw}(k)$, is:
@@
The $\beta$-Conditional value at risk ($\beta$-CVAR) is:
$\textrm{CVAR}_{\beta}^{zw}(k)=\frac{1}{1-\beta}\int_{L_{zw}(k,\delta)\geq\textrm{VAR}_{\beta}^{zw}(k)}L_{zw}(k,\delta)p(\delta)d\delta\;.$ (5)
In other words, when looking at the probability distribution of the loss function $L_{zw}(k,\delta)$, $\text{VAR}_{\beta}^{zw}(k)$ separates the $\beta$ most favorable cases from the $(1-\beta)$ worst cases, while $\textrm{CVAR}_{\beta}^{zw}(k)$ represents the conditional expected value of the loss associated with the worst cases $\delta$ relative to that loss being greater than $\text{VAR}_{\beta}^{zw}(k)$. This is schematically represented in Fig. 2.
Figure 2: Probability density $p^{zw}_{k}(\alpha)$ of the loss function $T_{zw}(\mathrm{s},k,\delta)||$ with $\beta$-VAR, $\beta$-CVAR and worst case
@@
which replaces problem (2), with the main differences: (i) the parameters $\delta$ are defined as random variables rather than deterministic parameters, over the same bounded support (Assumption 2.1), (ii) instead of minimizing the worst case without considering any probability, we seek to minimize the conditional expected value of the ($1-\beta_{i}$) worst cases.
3 Computation of VAR and CVAR
3.1 Formulation as an optimization problem
Definition 2.4 does not allow for a straightforward computation of the VAR and CVAR. Instead, one typically relies on the following result of Rockafellar [20], in which both are characterized as solutions of a convex optimization problem.
Proposition 3.1
(Expression of $\beta$-VAR and $\beta$-CVAR as solutions of a convex minimization problem [20])
Under Assumptions 2.1, 2.2 and 2.3, let us define the function
@@ 3.1 Formulation as an optimization problem
Corollary 3.2
Optimal controller for the $\beta$-CVAR problem
@@
3.2 Sample average approximation (SAA)
The function $F^{zw}_{\beta}(k,\alpha)$ can be approximated by sampling $N$ i.i.d. realizations $\tilde{\delta}^{1}$, $\tilde{\delta}^{2}$, … $\tilde{\delta}^{N}$ from the density $p$ and defining the sample average approximation (SAA) [15]
@@ 3.2 Sample average approximation (SAA)
where $A(k,\delta)$ is the dynamics matrix of the state-space representation of the closed loop, and $\lambda^{\mathcal{R}}_{\max}(\cdot)$ denotes the spectral abscissa. The constraint on the spectral abscissa ensures that Assumption 2.2 remains satisfied over the whole support $\mathcal{D}_{\delta}$ and not only over the samples $\tilde{\delta}^{(i)}$. How this constraint can be enforced is discussed in Section 4.1.
3.3 Confidence interval
To obtain an accurate solution to the original problem, it is necessary to choose $N$ sufficiently large. [10, Theorem 3] characterizes the convergence rate and derives the number of samples required to guarantee a prescribed accuracy level with a given confidence. However, this result also depends on the regularity of the uncertain loss function, which is not known a priori. In practice, thanks to the expression of VAR and CVAR as solutions of an optimization problem (Proposition 3.1), the procedure proposed in [17, Section 3.2] allows to assess the gap between the estimated $\beta$-CVAR:
@@ 3.3 Confidence interval
4 Optimization
4.1 Optimization procedure
@@
Initialization: Active configurations $\mathcal{D}_{\delta}^{a}=\{0\}$, number of samples $N$, tolerance $\epsilon$.
Step 1: Find $\hat{k}_{0}$ that robustly stabilizes $T_{zw}(s,k,\delta)$ over the whole support $\mathcal{D}_{\delta}$, and update $\mathcal{D}_{\delta}^{a}$.
Step 2: Form the SAAs with $N$ samples, initialize $k=\hat{k}_{0}$ and solve the SAA problem (12) where robust stability over the whole support:
@@
If $\lambda^{\mathcal{R}}_{\max}(A(\hat{k},\delta^{*}))>0$, add $\delta^{*}$ to $\mathcal{D}_{\delta}^{a}$ and go back to step 2. Otherwise, go to step 4.
@@
Step 4: Evaluate the CVARs of problem (12), with sufficiently tight intervals of confidence. If they are different from those obtained in step 2 by more than $\epsilon$%, increase $N$ and go back to step 2. Otherwise, terminate.
@@
Step 1 allows to provide an initial controller with robust stability over the whole support, as required by Assumption 2.2. It can be performed with [2], which relies on the algorithm of [4] and the active configurations approach described in next paragraph. In [4], the pure stabilization problem is addressed with the shifted $\mathcal{H}_{\infty}$ norm [9]. Optimization of the SAA approximation is performed in step 2; it can be addressed with deterministic nonsmooth optimization, cf. section 4.2. A worst-case search for unstable configurations is performed in step 3, as discussed in next paragraph. Finally, step 4 evaluates the estimated CVARs obtained with the optimized controller and their confidence intervals of Section 3.3 to validate or invalidate the controller.
The identification of the global worst-case stability configuration being NP-hard [29], the stability over the whole support $\mathcal{D}_{\delta}$ is addressed through the active configurations approach of [2]. It consists of enforcing stability of configurations belonging to a subset $\mathcal{D}_{\delta}^{a}\subset\mathcal{D}_{\delta}$, initialized with the nominal configuration, and updating this subset through (deterministic) worst-case searches (step 3). This procedure is necessary even when the initial controller stabilizes the system over the whole support (step 1), because this property may not be maintained when solving the SAA problem (step 2). If a destabilizing configuration $\delta^{*}$ is identified in step 3, it is added to $\mathcal{D}_{\delta}^{a}$, and step 2 is repeated. Note that the step 1 typically also uses this approach [2], and thus, it also updates $\mathcal{D}_{\delta}^{a}$.
4.2 Open challenges
@@
While we have proposed a general method to address the CVAR optimization, there are still open challenges to address the formulated problem in a systematic and efficient way.
@@
Future research should focus on developing suitable nonsmooth algorithms to address step 2 of Algorithm 1, efficiently and with convergence certificates. Of particular interest is the work of [18], which proposes an algorithm that tolerates inexact computation of the gradient or of the function itself. This algorithm is specifically designed for lower-$C^{1}$ functions, a class to which the SAA $\widetilde{F}^{zw}_{\beta}(k,\alpha)$ can easily be shown to belong under Assumptions 2.1, 2.2 and 2.3, since this property holds for $T_{zw}(s,k,\delta)||_{\infty}$ as a function of $k$ as long as $T_{zw}(s,k,\delta)$ is internally stable [2, Proposition 2]. In particular, [18] illustrates their method on the $\mathcal{H}_{\infty}$ norm where the exactness of the computation, based on a dichotomy involving the bounded real lemma, can be relaxed. This property is particularly desirable in our SAA where we have to evaluate many times the $\mathcal{H}_{\infty}$ norm, and where our SAA is itself an approximation.
@@
Finally, in this paper, the distributions are assumed to be perfectly known. When this assumption does not hold, distributionally robust control provides a natural extension; see, for example, [30, 14].
5 Application
5.1 Benchmark
We consider the benchmark of [16], namely the attitude control of a flexible spacecraft $G(s,\delta)$ to follow a reference $r$. The closed loop is represented in Fig. 3, where $d(s)$ represents actuator dynamics. The control input $u$ of the plant is the torque (3 components) applied to the satellite and the output vector contains the 3 angle measurements polluted by some noise $n$. We note $k\in\mathcal{K}$ the decision variables describing the controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, each one taking the measured angle and rate as inputs and generating a control torque. The closed-loop system has 50 states, 43 uncertain parameters $\delta_{i}$, and 72 decision parameters $k_{i}$.